Life expectancy Religion A Religion B Religion C No Religion
> 90 years old 27 26 18 29
Take α= 0.05.
H0 : PA = PB = PC = PNO RELIGION = .25
H1 : PA ≠ PB ≠ PC ≠ PNO RELIGION ≠ .25
a. What is the critical value of the chi-square?
b. What is the degree of freedom?
c. What is the value of the test statistics of the chi-square?
d. What is the p-value?
e. Can we conclude that the life expectancy of the Malaysian men depends on their religious affiliation? State your reason.
1
Expert's answer
2020-07-27T17:53:05-0400

a) using table critical value = 7.815

b)Degree of freedom=k-1= 4 - 1 = 3

c) X2= (27 -25)2/25+(26 -25)2/25+(18-25)2/25+(29 -25)2/25=2.8

d) using table - the P-Value is 0.4235.

e) The result is not significant at 0.05.  The life expectancy of the Malaysian men does not depend on their religious affiliation